# Tractatus Quanticum

### v1

@@ Abstract
What we hold onto is not knowledge of an absolute order, but the fleeting grasp of a perspective on perspectives. One that can look at others, and is hence aware of its own partiality. Good philosophy, seems to us, dismantles idols, rather than creating them, and, like a gentle friend, reminds us with indirect hints, over and over again, that the answer to the riddle is that there is no riddle.
@@ 1 Synthesis
    My knowledge about the world is a particular perspective.
@@ 2 Short
    My knowledge about the world is a particular [^1] Here, a ‘particular’ denotes a concrete instantiation of an assumed universal. The world is constituted by particular perspectives, while a universal, here, is just a sum of particulars. A sum to which it is always possible to add a new particular. perspective.
@@
      My knowledge of the world is the information I have or can have about the world.
@@
      Information has degrees of reliability. [^2] Cfr. De1989-FINPAC-2.
@@
      When the mutual information $I_{A:B}$ is maximal, knowing $b$ determines which $a$ is a fact, and $a$ is a fact relative to bb. [^3] RelativeinfoRovelli.
@@
      The self is the reflective point of a knowledge that includes information about itself. [^9] TIsmael2006-TISTSS
@@
      Within any perspective, circularity is broken indexically by that very perspective. [^10] TIsmael2006-TISTSS
@@ 3 Long
        The reality of which we speak can never be the total reality of all possible perspectives. [^11] Cfr. Laozi2003-LAOTDO, 1.1.
@@
          Reality consists in states of affairs, [^12] As expressed by Wittgensteintractatus proposition 2.01. but states of affairs do not exist outside a perspective.
@@
    My knowledge about the world is a particular [^13] Here, a ‘particular’ denotes a concrete instantiation of an assumed universal. The world is constituted by particular perspectives, while a universal, here, is just a sum of particulars. A sum to which it is always possible to add a new particular. perspective.
@@
      Information has degrees of reliability. [^14] Cfr. De1989-FINPAC-2.
@@
        Degrees of reliability can be quantified: $p(a)\in[0,1]$ quantifies my expectation that a possible fact $a$ is going to be the case if I acquire information about it. $p(a)=0$ indicates certainty that $a$ is the case, and $p(a)=1$ indicates certainty that $a$ is not the case.
@@
          A fact is what we call an element of physical reality. [^15] Cfr. Einstein1935-EINCQD, where an element of reality is defined in a very similar manner, as something I reliably expect to be true if observed.
@@
        The uncertainty in the probability $p(a)$ can be quantified by $H(p)=-\sum_{a}p(a)\ln p(a)$, which is called the Shannon’s entropy of p(a)p(a). [^16] Shannon
@@
        If I know which of $N$ alternatives is a fact, I have an information $I(N)=\ln N,$ Shannon’s first definition of amount of information. [^17] Shannon
@@
        A fuzzy logical space [^18] ZADEH1965338 can account for the situation when we are unable to predicate facts with $p(a)=1.$
@@
        By measuring $b$, the average increase in information about $a$, is given precisely by $I_{A:B}.$ [^19] RelativeinfoRovelli
@@
      When the mutual information $I_{A:B}$ is maximal, knowing $b$ determines which $a$ is a fact, and $a$ is a fact relative to bb. [^20] RelativeinfoRovelli
@@
        My knowledge of the world is a set of facts regarding the physical process I am, and correlations with the rest of the world. [^21] The emphasis on this point is novel with respect to Wittgensteintractatus. Knowledge and language are themselves physical facts, they can be also understood as such, and doing so clarifies.
@@
        In physical interactions and more in general in time, knowledge is acquired or lost.
@@
        Meaning is evolutionarily — or biologically — relevant information. [^22] Rovelli2018 and arcas2025intelligence.
@@
            These traces are not rare; they are not, like Dasein232323(Heidegger1962-HEIBAT-8). , specific of humans. They are generic effects of physical interaction upon systems.
@@
        The general form of dynamical laws is: if $a$ then $b$ with a certain reliability. That is, dynamical laws assign values to $p(b|a)$: likelihood of $b$ given $a$.
@@
          Physics is modal: it is not descriptive. It does not state or prescribes what is the case. It states what may be the case if something else is the case. It restricts the space of possibilities; it does not fix one single possibility.
@@
            The point 4.3.2.1.1. is what forbids us to take facts that we observe as absolute, beyond our perspective. These are the limits of naïve realism.
@@
        The facts that embody a perspective are represented by simultaneous values of variables. Therefore a perspective is defined by an abelian subalgebra of the algebra of variables. An abelian subalgebra is a linear subset of an algebra formed by elements that commute (such that AB=BAAB=BA). [^29] RelativeinfoRovelli.
@@
        The state of a system relative to a perspective is given by a linear functional $\rho$ (a “quantum state”) on the algebra of the variables of this system. Given the spectral decomposition $A=\sum_{a}a\,\Pi_{a}$, of a variable $A$, the probability $p(a)$ is given by $p(a)=\rho(\Pi_{a})$. For two commuting variables $A,B$, $p(a,b)=\rho(\Pi_{a}\Pi_{b})$. This shows that the quantum formalism includes the precise mathematical tools for defining relative facts, as detailed in 20.
@@
          We say that systems and in particular objects (which are bundles of properties303030(Simons1994-SIMPIP).) exist, insofar as they are manifest in a perspective (a relation). This is a decisive break from the older conception of things. [^31] See Wittgensteintractatus, prop. 3.25.
@@
        An atomic fact is the possibility of the value of a variable occurring in a perspective. [^32] See Wittgensteintractatus, prop. 2.0123.
@@
          It would then be impossible to draw up a picture of the world without a perspective. [^33] See Wittgensteintractatus, prop. 2.0212.
@@
        The maximum information available about a system is insufficient to determine the further information that the system can provide, because values of variables that do not commute with the known ones are not facts. Interactions with these variables bring about genuinely new information. This does not increase the maximum information about the system because it renders previous information irrelevant: some facts cease to be facts. [^37] The two postulates in axiomatic formulation of RQM, Rovelli1996-ROVRQM are (i) system’s relevant information is limited, but (ii) new information can always be acquired.
@@
      The self is the reflective point of a knowledge that includes information about itself. [^40] TIsmael2006-TISTSS
@@
        Knowledge about ourselves is always vastly incomplete.414141(Spinoza2020-SPISE).
@@
          Thought is a physical phenomenon; it can in principle be described as a set of facts.424242(Spinoza2020-SPISE). This is also what neurosciences aim to do. Ethics is a physical phenomenon; it can be also be in principle described as a set of facts. But they are both phenomena better described in appropriate layers as in 5.5.2..
@@
          The transparency of a perspective to another does not need to be realized at the physical level. It works more easily at higher levels, with the associated uncertainty. This is called “communicating”. [^43] Zhuangzi, 17.
@@
          There is no un-embodied language. [^44] This is the the central difference from Wittgensteintractatus.
@@
        These same variables can be accounted for from another perspective. But there is no way to account for all perspectives, because of $1.4.2,3.10,$ and $5.4.$ A world is only such with respect to something physical, when quantum phenomena are not disregarded. This gives rise to the circularity of making sense of perspective, from other perspectives, as correlations.
@@
          On logical analysis, circularity is already employed in set theory through the theorem stating that “every graph has a unique decoration.” [^45] Aczel1988-ACZNS chapter $6$. In this way, it becomes possible to define a graph in terms of itself, something forbidden in Zermelo–Frankel set theory. This is useful circularity.
@@
      3. 6.8.3.
        The idea of the ensemble formed by the totality of all perspectives, and of the facts within them, does not belong to reality, for it is not contained within any perspective. [^47] The emphasis on this point is connected to Wittgensteintractatus (Proposition 5.6), where the author reflects on the limits of language: the boundary between what can be said and what can only be shown.
        1. 6.8.3.1.
@@
      1. 6.9.1.
        Hence all of the above is just a perspective on perspectives, and as such partial. It is not a fundamental view.484848(Ngrjuna) $MMK.13.8$.
      2. 6.9.2.

### v2

# Tractatus Quanticus

@@ Abstract
What we hold onto is not knowledge of an absolute order, but the fleeting grasp of a perspective on perspectives. Hence this is at most just a manifesto for a picture of the world suggested by quantum mechanics. Each proposition attempts to show a possible way of formulating questions about nature in a manner coherent with what we have recently learned about it. A manner aware of its own partiality. Good philosophy, seems to us, dismantles idols, rather than creating them, and, like a gentle friend, reminds us with indirect hints, over and over again, that the answer to the riddle is that there is no riddle.
@@
    My knowledge about the world is a particular kind of perspective.
@@
    My knowledge about the world is a particular kind of perspective. [^1] Here, “particular kind” denotes a concrete instantiation of what is assumed to be universal. The world is constituted by particular perspectives, whereas a universal is merely the sum of these particulars. A whole that can always be extended by adding a further particular.
@@
      My knowledge [^2] We use the term “knowledge” in the Wittgenstein (1953)’s sense §116, where knowing is embedded in practices and does not imply access to a total view or a metaphysical picture. of the world is the information I have or can have about the world.
@@
      Information has degrees of reliability. [^3] Cfr. De Finetti (1989).
@@
      When the relative information $I_{A|B}$ is maximal, knowing $b$ determines which $a$ is a fact, and $a$ is a fact relative to bb. [^4] Di Biagio and Rovelli (2025).
@@
      Physics is modal: the general form of dynamical laws is: if $a$ then $b$ with a certain reliability. Dynamical laws assign values to conditional probabilities $p(b|a)$: likelihood of $b$ given $a$.
    5. 4.5.
      Causation is the aspect of modality that can support intervention.
    6. 4.6.
      My world is coloured by my interaction with it. [^5] Price (2007).
    7. 4.7.
@@
      The self is the reflective point of a knowledge that includes information about itself. [^11] Ismael (2006)
@@
      Physics itself is in a perspective.
@@
      Within any perspective, circularity is broken indexically by that very perspective. [^12] Ismael (2006)
    6. $6.$
@@
      1. 1.4.1.
        At this point, an attentive reader might be led to question the absoluteness of physical rules; this issue will be addressed in 6.6.4.,as we must climb the ladder one rung at a time in order to fully grasp the lesson proposed by physics.
@@
        The reality of which we speak can never be the total reality of all possible perspectives. [^13] Cfr. Laozi and Ivanhoe (2003), 1.1.
@@
          Reality consists in states of affairs, [^14] As expressed by Wittgenstein (1922) proposition 2.01. but states of affairs do not exist outside a perspective.
@@
    My knowledge about the world is a particular kind of perspective.
@@
      Information has degrees of reliability. [^15] Cfr. De Finetti (1989).
@@
        Degrees of reliability can be quantified: $p(a)\in[0,1]$ quantifies my expectation that a possible fact $a$ is going to be the case if I acquire information about it. $p(a)=1$ indicates certainty that $a$ is the case, and $p(a)=0$ indicates certainty that $a$ is not the case.
@@
          A fact is what we call an element of physical reality. [^16] Cfr. Einstein et al. (1935), where an element of reality is defined in a very similar manner, as something I reliably expect to be true if observed.
@@
        The uncertainty in the probability $p(a)$ can be quantified by $H(p)=-\sum_{a}p(a)\ln p(a)$, which is called the Shannon’s entropy of p(a)p(a). [^17] Shannon (1948)
@@
        If I know which of $N$ alternatives is a fact, I have an information $I(N)=\ln N,$ Shannon’s first definition of amount of information. [^18] Shannon (1948)
@@
        A fuzzy logical space [^19] Zadeh (1965) can account for the situation when we are unable to predicate facts with $p(a)=1.$
@@
        By measuring $b$, the average increase in information about $a$, is given precisely by $I_{A:B}.$ [^20] Di Biagio and Rovelli (2025)
@@
      When the relative information $I_{A|B}$ is maximal, knowing $b$ determines which $a$ is a fact, and $a$ is a fact relative to bb. [^21] Di Biagio and Rovelli (2025)
@@
        My knowledge of the world is a set of facts regarding the physical process I am, and correlations with the rest of the world. [^22] The emphasis on this point is novel with respect to Wittgenstein (1922). Knowledge and language are themselves physical facts, they can be also understood as such, and doing so clarifies.
@@
        In physical interactions and more in general in time, knowledge either is acquired or lost.
@@
        In the classical limit, the concept of information acquires an epistemic dimension: the observer’s –possibly partial– knowledge about the objective perspective-independent state of affairs.
@@
          The sharp distinction between facts and information about them makes only sense assuming absolute states of facts (see 4.4.8.4.4.8.4.3.).
        2. 3.10.3.2.
          In the quantum world, mutual information is always a measure of physical correlations.
          1. 3.10.3.2.1.
            The von Neumann’s entropy of a density matrix $\rho$ quantifies physical correlation.
            1. 3.10.3.2.1.1.
              By contrast, the measure introduced in $3.10.3.2$ quantifies entanglement correlations between subsystems, rather than epistemic knowledge.
        3. 3.10.3.3.
          We can recover of an epistemic notion of information in the sense of comparing –say– limited information on a system available in a perspective, with richer information about the same system available in –say– our own perspective.
          1. 3.10.3.3.1.
            For example when we quantify our knowledge of a quantum system, through the von Neumann entropy of the global state, we are adopting once again an epistemic notion of information.
      4. 3.10.4.
        Meaning is evolutionarily — or biologically — relevant information. [^23] Rovelli (2018) and Arcas (2025).
        1. 3.10.4.1.
@@
          1. 3.10.4.3.1.
            Information is not a primitive ontological ingredient; it is a structural feature of the correlations between systems. [^24] This marks a departure from the information-theoretic interpretation proposed by Bub and Pitowsky (2010).
@@
            These traces are not rare; they are not, like Dasein252525(Heidegger, 1962). , specific of humans. They are generic effects of physical interaction upon systems.
@@
        Physics is modal: the general form of dynamical laws is: if $a$ then $b$ with a certain reliability. That is, dynamical laws assign values to $p(b|a)$: likelihood of $b$ given $a$.
@@
          Physics is not descriptive. It does not state or prescribes what is the case. It states what may be the case if something else is the case. It restricts the space of possibilities; it does not fix one single possibility.
@@
            The point 4.3.2.2. is what forbids us to take facts that we observe as absolute, beyond our perspective. These are the limits of naïve realism.
@@
        The facts that embody a perspective are represented by simultaneous values of variables. Therefore a perspective is defined by an abelian subalgebra of the algebra of variables. An abelian subalgebra is a linear subset of an algebra formed by elements that commute (such that AB=BAAB=BA). [^31] Di Biagio and Rovelli (2025).
@@
        The state of a system relative to a perspective is given by a linear functional $\rho$ (a “quantum state”) on the algebra of the variables of this system. Given the spectral decomposition $A=\sum_{a}a\,\Pi_{a}$, of a variable $A$, the probability $p(a)$ is given by $p(a)=\rho(\Pi_{a})$. For two commuting variables $A,B$, $p(a,b)=\rho(\Pi_{a}\Pi_{b})$. This shows that the quantum formalism includes the precise mathematical tools for defining relative facts, as detailed in 21.
@@
          We say that systems and in particular objects (which are bundles of properties323232(Simons, 1994).) exist, insofar as they are manifest in a perspective (a relation). This is a decisive break from the older conception of things. [^33] See Wittgenstein (1922), §3.25.
@@
        An atomic fact is the possibility of the value of a variable occurring in a perspective. [^34] See Wittgenstein (1922), §2.0123.
@@
          It would then be impossible to draw up a picture of the world without a perspective . [^35] See Wittgenstein (1922), §2.0212.
@@
        The maximum information available about a system is insufficient to determine the further information that the system can provide, because values of variables that do not commute with the known ones are not facts. Interactions with these variables bring about genuinely new information. This does not increase the maximum information about the system because it renders previous information irrelevant: some facts cease to be facts. [^39] The two postulates in axiomatic formulation of RQM, Rovelli (1996) are (i) system’s relevant information is limited, but (ii) new information can always be acquired.
@@
      1. 5.1.1.
        What I know or can know about the events $a$ and $b$ is given by their joint probability $p(a,b)$. From this, the perspective of $b$ on $a$ is defined by the conditional probabilities $p(a|b)$.
      2. 5.1.2.
        Knowledge, as stated in 3.3.10. is physically underpinned by a physical configuration correlated with the part of the world it concerns. Because it is physical, it is accessible from other perspectives through interaction. No perspective lies beyond the physical.
@@
      The self is the reflective point of a knowledge that includes information about itself. [^42] Ismael (2006)
@@
        Knowledge about ourselves is always vastly incomplete.434343(Spinoza, 2020).
@@
          Thought is a physical phenomenon; it can in principle be described as a set of facts.444444(Spinoza, 2020). This is also what neurosciences aim to do. Ethics is a physical phenomenon; it can be also be in principle described as a set of facts. But they are both phenomena better described in appropriate layers as in 5.5.2..
@@
          The transparency of a perspective to another does not need to be realized at the physical level. It works more easily at higher levels, with the associated uncertainty. This is called “communicating”. [^45] Watson (2003), 17.
@@
          There is no un-embodied language. [^46] This is the the central difference from Wittgenstein (1922).
@@
        These same variables can be accounted for from another perspective. But there is no way to account for all perspectives, because of $1.4.2$, 3.3.10. and $5.4.$ A world is only such with respect to something physical, when quantum phenomena are not disregarded. This gives rise to the circularity of making sense of perspective, from other perspectives, as correlations.
@@
          On logical analysis, circularity is already employed in set theory through the theorem stating that “every graph has a unique decoration.” [^47] Aczel (1988) chapter $6$. In this way, it becomes possible to define a graph in terms of itself, something forbidden in Zermelo–Frankel set theory. This is useful circularity.
@@
      Physics itself is in a perspective.
      1. 6.4.1.
        In physics, we use variables, functions and rules, and mistakenly consider these absolute. This gives rise to the wrong idea of a series of absolute rules outside any perspective. But the rules, in fact, are part of our knowledge, and as such perspectival themselves.
    5. 6.5.
@@
      3. 6.9.3.
        The idea of the ensemble formed by the totality of all perspectives, and of the facts within them, does not belong to reality, for it is not contained within any perspective. [^49] The emphasis on this point is connected to Wittgenstein (1922) (Proposition 5.6), where the author reflects on the limits of language: the boundary between what can be said and what can only be shown.
        1. 6.9.3.1.
@@
      1. 6.10.1.
        Hence all of the above is just a perspective on perspectives, and as such partial. It is not a fundamental view.505050(Siderits, 2022) $MMK.13.8$.
      2. 6.10.2.